OpenStudy (anonymous):

PLEASE HELP .. On January 1 2010, Chessville has a population growth. Its population increases 7% each year. On the same day checkersville has a population of 70,000 people. Cheakersveille starts to expeirence a population decline . It's population decreases 4% each year . During what year will the population of Chessville first exceed that of Checkersville?

OpenStudy (anonymous):

we are missing the initial population of "chessville"

OpenStudy (anonymous):

oh sorry its 50000

OpenStudy (anonymous):

\[50000(1.07)^t=70,000(.96)^t\] solve for \(t\)

OpenStudy (anonymous):

lets see, divide by \(50,000\) get \[(1.07)^t=1.4(.96)^t\] divide by \((.96)^t\) get \[\left(\frac{1.07}{.96}\right)^t=1.4\]

OpenStudy (anonymous):

then use "change of base" formula to find \(t\) \[t=\frac{\ln(1.4)}{\ln(\frac{1.07}{.96})}\]

OpenStudy (anonymous):

and of course a calculator